# Does the 220 PeV Event at KM3NeT Point to New Physics?

### v1

@@ Abstract
In this work, we discuss two scenarios that address this tension. What they have in common is a sterile neutrino. We assume that sterile neutrinos get copiously produced at the source and dominate in flux over active neutrinos. If that were not the case, active neutrino fluxes at the KM3NeT and IceCube detectors would be similar, given the propagation distances in Earth which do not facilitate attenuation of neutrino flux for Standard Model (SM) cross sections. Even if there was a significant attenuation, this would reduce the flux at KM3NeT relative to IceCube, which would further strengthen the tension between the two experiments and is opposite from what we want to achieve. Sterile neutrinos could be produced directly from dark matter annihilation or decay Morrison:2022zwk ; Brdar:2016thq and if they are sufficiently light they will not decay to SM particles before reaching Earth, irrespective of the coupling between SM and the sterile neutrino sector. Another possibility is that the point source is very dense and only weakly interacting particles, such as sterile neutrinos, can exit the production region with significant energy. The lack of multi-messenger observations KM3NeT:2025bxl in association to KM3-230213A event further supports our premise.
@@
The KM3NeT collaboration has very accurately determined the direction of the incoming $\mathcal{O}(100)$ PeV neutrino trajectory, with an uncertainty of $1.5^{\circ}$. The origin of KM3-230213A event corresponds to an incidence angle of $0.6^{\circ}$ at KM3NeT and $\sim 8^{\circ}$ above the horizon at IceCube. Using this, it is straightforward to determine the path length through rock and sea for KM3NeT, as well as the respective path length through ice en route to the IceCube detector for neutrinos coming from the same source. The topography is illustrated in Fig. 1, where we show that the path through Earth for KM3NeT is an order of magnitude longer than that for IceCube. We employ this difference in propagation distance to alleviate the tension between IceCube and KM3NeT. Namely, in what follows, we discuss two mechanisms where sterile neutrino flux partially transitions into active neutrino flux inside the Earth. Due to the longer propagation distance inside the Earth, the active neutrino flux at KM3NeT will be significantly higher compared to IceCube, compensating for the smaller effective area and shorter data-taking time, thereby alleviating the tension.
@@
Neutrino production through matter-induced resonance. When the vacuum mixing between active and sterile neutrinos is very small, matter can resonantly amplify sterile-to-active transitions. The MSW resonance Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj is realized when the matter potential difference for the involved neutrino species, $V$, equals $[\Delta m^{2}/(2E_{\nu})]\cos 2\theta$, where $\Delta m^{2}$ and $\theta$ are the mass-squared difference and vacuum mixing angle, respectively. Taking the values from the SM, matter potential reads $V\equiv V_{\text{CC}}=\sqrt{2}G_{F}n_{e}\simeq 3.7\times 10^{-23}$ GeV for matter with a density of $1\,\text{g}/\text{cm}^{3}$. The resonant oscillation length then reads $L_{\text{res}}\simeq\pi/(V\sin 2\theta)\approx 5\times 10^{3}\,\theta^{-1}$ km in the limit of small $\theta$. For keV-scale sterile neutrino, which will turn out to be required in our analysis, maximum mixing angle value of $\theta^{2}\simeq 10^{-3}$ is allowed in light of terrestrial experiments MINOS:2017cae ; Bolton:2019pcu (such values are in tension with cosmological constraints, but those can be alleviated Farzan:2019yvo ; Cherry:2016jol ; Chu:2018gxk ). This yields $L_{\text{res}}\simeq 10^{5}$ km or longer, and for $L\approx 150$ km, which corresponds to KM3NeT, the oscillation probability would then not surpass $L^{2}/L_{\text{res}}^{2}\approx 10^{-6}$.
@@
This expression is obtained after integrating out $V^{\mu}$ (note that the mass of the introduced vector boson, $m_{V}$, always satisfies $m_{V}^{2}\gg Q^{2}=0$). In Eq. 2, $p$ and $n$ represent protons and neutrons, and $G_{B}=g_{b}g_{b}^{\prime}/m_{V}^{2}$. The sterile neutrino matter potential can then be written as $V_{s}=G_{B}n_{\text{nuc}}$ Pospelov:2011ha ; Kopp:2014fha , where $n_{\text{nuc}}=n_{p}+n_{n}$ is the nucleon number density in matter. For $V_{s}\gg G_{F}n_{e}$, the difference between the sterile and active neutrino matter potential, $V$, is approximately equal to $V_{s}$. We parameterize $V_{s}$ using the dimensionless quantity $\epsilon_{ss}=G_{B}/(\sqrt{2}G_{F})$, which represents the ratio of the new baryonic interaction strength to Fermi’s interaction strength. In the literature Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm , $\epsilon_{ss}$ values up to $\mathcal{O}(10^{2}-10^{3})$ have been explored. Upon investigating the effects of $\epsilon_{ss}$ on neutrino oscillations, we find that values $|\epsilon_{ss}|<10^{3}$ typically lead to deviations smaller than a few percent for atmospheric neutrinos. Note that such an interaction strength precisely addresses the $L^{2}/L_{\text{res}}^{2}\ll 1$ problem for KM3NeT and makes the propagation length in Earth comparable to the resonant oscillation length. We should stress that limits on $\epsilon_{ss}$ become more severe if a sterile neutrino mixes with several active neutrinos; constraints arising from solar neutrino data in a scenario where a sterile neutrino mixes with an electron neutrino are particularly restrictive Kopp:2014fha . However, in this work, we consider sterile neutrino mixing only with a muon neutrino. The generation of muon neutrino flux via oscillations is crucial, as muon neutrinos can produce muon tracks at KM3NeT through SM charged current interactions, explaining the recent observation. For $E_{\nu}=220$ PeV and propagation in water/rock ($\rho\approx 1-2.2\,\text{g}/\text{cm}^{3}$, $n_{\text{nuc}}\approx\rho/m_{p}$), MSW resonance occurs for $\sqrt{\Delta m^{2}}\approx m_{s}\simeq 2\times 10^{-1}\sqrt{\epsilon_{ss}}$ keV, where $m_{s}$ is the sterile neutrino mass. For $\epsilon_{ss}\simeq\mathcal{O}(100)$, a keV-scale sterile neutrino will undergo resonance in the Earth en route to the KM3NeT detector. We do not assume constant density matter, as propagation of sterile neutrinos in the direction corresponding to the detected event at KM3NeT involves passage through both rock and water (see Fig. 1); we take the former (latter) to be $100$ km ($47$ km) long. To this end, we solve the Schrödinger equation which is governed by the following Hamiltonian in the ($\nu_{\mu}$, $\nu_{s}$) flavor basis
$H=\!\begin{pmatrix}c_{\theta}&s_{\theta}\\ -s_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}0&0\\ 0&\frac{m_{s}^{2}}{2E_{\nu}}\end{pmatrix}\begin{pmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\end{pmatrix}\!+\!\begin{pmatrix}V_{\text{NC}}&0\\ 0&V_{s}\end{pmatrix},$ (3)
@@
where we have abbreviated $c_{\theta}=\cos\theta$ and $s_{\theta}=\sin\theta$. The term $V_{\text{NC}}=-G_{F}n_{n}/\sqrt{2}\approx-V_{\text{CC}}/2$ is the SM neutral current matter potential which, while included in our analysis, will not impact results due to its smallness with respect to $V_{s}=\sqrt{2}G_{F}\epsilon_{ss}(n_{n}+n_{p})$. The resonance occurs in the neutrino sector for negative values of $\epsilon_{ss}$ and this is a scenario that we will adopt (for positive values of $\epsilon_{ss}$ resonance would be observed in the antineutrino sector).
@@
Figure 2: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown in the parameter space of the sterile neutrino mass $m_{s}$ and interaction strength $-\epsilon_{ss}$ for both KM3NeT and IceCube. The vacuum mixing angle is fixed to $\theta=10^{-2}$. Deeper shades of blue and red correspond to higher oscillation probabilities at KM3NeT and IceCube, respectively.
@@
Figure 3: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown as a function of interaction strength $-\epsilon_{ss}$ for both KM3NeT and IceCube, for a sterile neutrino mass of $m_{s}=3$ keV and a vacuum mixing angle $\theta=10^{-2}$.
@@
The results of flavor evolution in the Earth for both IceCube and KM3NeT are shown in Figs. 2 and 3; in both figures, the vacuum mixing angle is fixed to $\theta=10^{-2}$ and we present oscillation probabilities after traversing 147 km and 14 km for KM3NeT and IceCube, respectively. In Fig. 2, we show sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the parameter space of sterile neutrino mass (scanned between $1$ and $10$ keV, values for which a resonance is expected) and $-\epsilon_{ss}$. For KM3NeT, sizable values of $P_{s\mu}$ already appear for $-\epsilon_{ss}\gtrsim 30$. In the region between $-\epsilon_{ss}\simeq 100$ and $-\epsilon_{ss}\simeq 10^{3}$, $P_{s\mu}$ for KM3NeT attains values of $P_{s\mu}\gtrsim 10^{-2}$, implying a relatively large muon neutrino flux generated in Earth. At around $-\epsilon_{ss}\sim 10^{3}$, the oscillation probability for IceCube reaches $P_{s\mu}\simeq 10^{-3}$, remaining smaller by 11112222 orders of magnitude compared to that of KM3NeT. This is because $P_{s\mu}$ depends on $\sin^{2}(L/L_{\text{res}})$, and the neutrino propagation length in Earth (along the direction of the 220 PeV neutrino detected by KM3NeT) is about 10 times shorter at IceCube. Therefore, at resonance, $P_{s\mu}$ is always expected to be at least two orders of magnitude larger for KM3NeT. The figure can be extended to larger values of $-\epsilon_{ss}$ that would require heavier sterile neutrinos to induce a resonance. However, increasing the interaction strength would create tension with the data Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm . In Fig. 3, we show $P_{s\mu}$ as a function of $-\epsilon_{ss}$ for both KM3NeT and IceCube and for a fixed sterile neutrino mass of $m_{s}=3$ keV. The resonance is obtained for $-\epsilon_{ss}\simeq 150$, a parameter point where $P_{s\mu}$ for KM3NeT is two orders of magnitude larger than for IceCube. This difference in oscillation probability, i.e., in the generated muon neutrino flux, alleviates the tension between the two experiments by compensating for KM3NeT’s smaller effective area and shorter data-taking period. In Fig. 3, we also show the dotted line, which represents the probability for sterile-to-active oscillations in vacuum, i.e. the muon neutrino flux generated between the source and the Earth relative to the sterile neutrino flux. The muon neutrino flux generated in vacuum is subdominant compared to the one generated through resonance, as required, since otherwise the event rate would be dominated by neutrinos produced outside the Earth, which would not explain the tension.
@@
Figure 4: The sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the scenario with off-diagonal non-standard neutrino interactions is shown for both KM3NeT and IceCube for $m_{s}=100$ eV.
In our scenario, a sterile neutrino with $E_{\nu}=220$ PeV would mix with the active neutrino sector due to matter effects parameterized by $\epsilon_{\mu s}$. The effective mixing angle in matter in the limit of a tiny vacuum mixing angle $\theta$ and large $m_{s}$ takes the form
@@
In the limit of $V_{\text{CC}}\gg m_{s}^{2}/E_{\nu}$, this further reduces to $\theta_{m}=(1/2)\tan^{-1}(4\epsilon_{\mu s})$. The effective mass squared difference in matter for $\theta\approx 0$ reads
@@
where $L$ is neutrino propagation length in Earth. In Fig. 4, we show the conversion probabilities for KM3NeT and IceCube as a function of $\epsilon_{\mu s}$ for $m_{s}=100\text{ eV}$ and $\theta=10^{-4}$. We have scrutinized the impact of the considered interaction on $P_{ee}$, $P_{\mu e}$, and $P_{\mu\mu}$ at low energies and found that $\epsilon_{\mu s}$ values up to $\mathcal{O}(1-10)$ have negligible effect on neutrino oscillations for $m_{s}\gtrsim 100$ eV. Hence, in the figure, we present oscillation probabilities for $\epsilon_{\mu s}$ values up to $10$. The sterile-to-active oscillation probability for KM3NeT is larger than that for IceCube by roughly two orders of magnitude for all presented values of $\epsilon_{\mu s}$. This can be understood from the oscillation formula in Eq. 8, which features an $L^{2}$ dependence when ${m_{\text{eff}}^{2}L}/{(4E_{\nu})}\lesssim\mathcal{O}(1)$, which is satisfied. Thus, the difference in the propagation length between KM3NeT and IceCube leads to the observed difference in the value of $P_{s\mu}$. While the ratio of $P_{s\mu}$ for KM3NeT and IceCube does not depend on $\epsilon_{\mu s}$, it should be noted that solutions with larger $\epsilon_{\mu s}$ are more favorable, as they lead to a larger generated active neutrino flux at KM3NeT, and thus require smaller incoming flux of sterile neutrinos. As in our previous scenario with resonance, the larger muon neutrino flux generated en route to KM3NeT alleviates the tension arising from the non-observation of $\mathcal{O}(100)$ PeV neutrinos in IceCube.

### v2

@@ Abstract
In this work, we discuss two scenarios that address this tension. What they have in common is a sterile neutrino. We assume that sterile neutrinos get copiously produced at a transient source and dominate in flux over active neutrinos. If that were not the case, active neutrino fluxes at the KM3NeT and IceCube detectors would be similar, given the propagation distances in Earth which do not facilitate attenuation of neutrino flux for Standard Model (SM) cross sections. Even if there was a significant attenuation, this would reduce the flux at KM3NeT relative to IceCube, which would further strengthen the tension between the two experiments and is opposite from what we want to achieve. High-energy sterile neutrinos could be produced directly in a transient source that is extremely dense. Sources such as choked GRBs or AGNs with a large column density of matter directly in the line of sight are promising candidates, where only weakly interacting particles, such as sterile neutrinos, could exit the production region with significant energy. Alternatively, dark sector interaction that stops active neutrinos could also be invoked, where only the sterile counterpart could traverse through a dense layer of dark sector and astrophysical distances. The lack of multi-messenger observations KM3NeT:2025bxl ; Fang:2025nzg in association to KM3-230213A further supports our premise.
@@
The KM3NeT collaboration has very accurately determined the direction of the incoming $\mathcal{O}(100)$ PeV neutrino trajectory, with an uncertainty of $1.5^{\circ}$. The origin of KM3-230213A event corresponds to an incidence angle of $0.6^{\circ}$ at KM3NeT and $\sim 8^{\circ}$ above the horizon at IceCube. Using this, it is straightforward to determine the path length through rock and sea for KM3NeT, as well as the respective path length through ice en route to the IceCube detector for neutrinos coming from the same source. The topography is illustrated in Fig. 1, where we show that the path through Earth for KM3NeT is an order of magnitude longer than that for IceCube. We employ this difference in propagation distance to alleviate the tension between IceCube and KM3NeT. Note that, a single transient source is essential, as it uniquely fixes the Earth-traversed chord length and prevents IceCube from sampling comparable path-lengths for similar events happening at different coordinates of the sky. However, this particular source can be as long-lived as $\mathcal{O}(1)$ year, as the location of IceCube at the South Pole means that the path length traversed through Earth would constantly be $\mathcal{O}(10)$ km, whereas for KM3NeT, Earth’s rotation would allow it to have the optimal Line-of-Sight distance through Earth matter during some time of the day. Namely, in what follows, we discuss two mechanisms where sterile neutrino flux partially transitions into active neutrino flux inside the Earth. Due to the longer propagation distance inside the Earth, the active neutrino flux at KM3NeT will be significantly higher compared to IceCube, compensating for the smaller effective area, and thereby alleviating the tension.
@@
Neutrino production through matter-induced resonance. When the vacuum mixing between active and sterile neutrinos is very small, matter can resonantly amplify sterile-to-active transitions. The MSW resonance Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj is realized when the matter potential difference for the involved neutrino species, $V$, equals $\Delta m^{2}/(2E_{\nu})]\cos 2\theta$, where $\Delta m^{2}$ and $\theta$ are the mass-squared difference and vacuum mixing angle, respectively. Taking the values from the SM, matter potential reads $V\equiv V_{\text{CC}}=\sqrt{2}G_{F}n_{e}\simeq 3.7\times 10^{-23}$ GeV for matter with a density of $1\,\text{g}/\text{cm}^{3}$. The resonant oscillation length then reads $L_{\text{res}}\simeq\pi/(V\sin 2\theta)\approx 5\times 10^{3}\,\theta^{-1}$ km in the limit of small $\theta$. For keV-scale sterile neutrino, which will turn out to be required in our analysis, maximum mixing angle value of $\theta^{2}\simeq 10^{-3}$ is allowed in light of terrestrial experiments MINOS:2017cae ; Bolton:2019pcu . Such values are in tension with cosmological constraints, but those can be alleviated Dasgupta:2013zpn ; Cherry:2016jol ; Chu:2018gxk ; Farzan:2019yvo ; Cline:2019seo ; Gelmini:2019esj ; Barbieri:2025moq . This yields $L_{\text{res}}\simeq 10^{5}$ km or longer, and for $L\approx 150$ km, which corresponds to KM3NeT, the oscillation probability would then not surpass $L^{2}/L_{\text{res}}^{2}\approx 10^{-6}$.
@@
This expression is obtained after integrating out $V^{\mu}$ (note that the mass of the introduced vector boson, $m_{V}$, always satisfies $m_{V}^{2}\gg Q^{2}=0$). In Eq. 2, $p$ and $n$ represent protons and neutrons, and $G_{B}=g_{b}g_{b}^{\prime}/m_{V}^{2}$. The sterile neutrino matter potential can then be written as $V_{s}=G_{B}n_{\text{nuc}}$ Pospelov:2011ha ; Kopp:2014fha , where $n_{\text{nuc}}=n_{p}+n_{n}$ is the nucleon number density in matter. For $V_{s}\gg G_{F}n_{e}$, the difference between the sterile and active neutrino matter potential, $V$, is approximately equal to $V_{s}$. We parameterize $V_{s}$ using the dimensionless quantity $\epsilon_{ss}=G_{B}/(\sqrt{2}G_{F})$, which represents the ratio of the new baryonic interaction strength to Fermi’s interaction strength. In the literature Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm , $\epsilon_{ss}$ values up to $\mathcal{O}(10^{2}-10^{3})$ have been explored. Upon investigating the effects of $\epsilon_{ss}$ on neutrino oscillations, we find that values $|\epsilon_{ss}|<10^{3}$ typically lead to deviations smaller than a few percent for atmospheric neutrinos. Note that such an interaction strength precisely addresses the $L^{2}/L_{\text{res}}^{2}\ll 1$ problem for KM3NeT and makes the propagation length in Earth comparable to the resonant oscillation length. We should stress that limits on $\epsilon_{ss}$ become more severe if a sterile neutrino mixes with several active neutrinos; constraints arising from solar neutrino data in a scenario where a sterile neutrino mixes with an electron neutrino are particularly restrictive Kopp:2014fha . However, in this work, we consider sterile neutrino mixing only with a muon neutrino. The generation of muon neutrino flux via oscillations is crucial, as muon neutrinos can produce muon tracks at KM3NeT through SM charged current interactions, explaining the recent observation. For $E_{\nu}=220$ PeV and propagation in water/rock ($\rho\approx 1-2.2\,\text{g}/\text{cm}^{3}$, $n_{\text{nuc}}\approx\rho/m_{p}$), MSW resonance occurs for $\sqrt{\Delta m^{2}}\approx m_{s}\simeq 2\times 10^{-1}\sqrt{|\epsilon_{ss}|}$ keV, where $m_{s}$ is the sterile neutrino mass. For $\epsilon_{ss}\simeq\mathcal{O}(100)$, a keV-scale sterile neutrino will undergo resonance in the Earth en route to the KM3NeT detector. Propagation of sterile neutrinos in the direction corresponding to the detected event at KM3NeT involves passage through both rock and water (see Fig. 1); we take the former (latter) to be $100$ km ($47$ km) long, with rock (water) density of $2.2~\rm{g}/\rm{cm}^{3}$ ($1~\rm{g}/\rm{cm}^{3}$). We assume constant density matter, with an average density of $1.82~\rm{g}/\rm{cm}^{3}$. A more detailed treatment with varying densities would only change this picture marginally. To this end, we solve the Schrödinger equation which is governed by the following Hamiltonian in the ($\nu_{\mu}$, $\nu_{s}$) flavor basis
$\hskip-3.0ptH=\!\begin{pmatrix}c_{\theta}&s_{\theta}\\ -s_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}0&0\\ 0&\frac{m_{s}^{2}}{2E_{\nu}}\end{pmatrix}\begin{pmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\end{pmatrix}\!+\!\begin{pmatrix}V_{\text{NC}}&0\\ 0&V_{s}\end{pmatrix}$, (3)
@@
where we have abbreviated $c_{\theta}=\cos\theta$ and $s_{\theta}=\sin\theta$. The term $V_{\text{NC}}=-G_{F}n_{n}/\sqrt{2}\approx-V_{\text{CC}}/2$ is the SM neutral current matter potential which, while included in our analysis, will not impact results due to its smallness with respect to $V_{s}=\sqrt{2}G_{F}\epsilon_{ss}(n_{n}+n_{p})$. Further, we simplify $n_{p}\approx n_{n}$, i.e., $V_{s}\approx 2\epsilon_{ss}V_{\rm CC}$. The resonance occurs in the neutrino sector for negative values of $\epsilon_{ss}$ and this is a scenario that we will adopt (for positive values of $\epsilon_{ss}$ resonance would be observed in the antineutrino sector).
@@
Figure 2: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown in the parameter space of the sterile neutrino mass $m_{s}$ and interaction strength $\epsilon_{ss}$ for both KM3NeT and IceCube. The vacuum mixing angle is fixed to $\theta=10^{-2}$. Deeper shades of blue and red correspond to higher oscillation probabilities at KM3NeT and IceCube, respectively.
@@
Figure 3: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown as a function of interaction strength $\epsilon_{ss}$ for both KM3NeT and IceCube, for a sterile neutrino mass of $m_{s}=3$ keV and a vacuum mixing angle $\theta=10^{-2}$.
@@
The results of flavor evolution in the Earth for both IceCube and KM3NeT are shown in Figs. 2 and 3 (for more details regarding the energy dependence, see the Supplemental Material supplemental ); in both figures, the vacuum mixing angle is fixed to $\theta=10^{-2}$ and we present oscillation probabilities after traversing 147 km and 14 km for KM3NeT and IceCube, respectively. In Fig. 2, we show sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the parameter space of sterile neutrino mass (scanned between $1$ and $10$ keV, values for which a resonance is expected) and $\epsilon_{ss}$. For KM3NeT, sizable values of $P_{s\mu}$ already appear for $\epsilon_{ss}\gtrsim 30$. In the region between $\epsilon_{ss}\simeq 100$ and $\epsilon_{ss}\simeq 10^{3}$, $P_{s\mu}$ for KM3NeT attains values of $P_{s\mu}\gtrsim 10^{-2}$, implying a relatively large muon neutrino flux generated in Earth. At around $\epsilon_{ss}\sim 10^{3}$, the oscillation probability for IceCube reaches $P_{s\mu}\simeq 10^{-3}$, remaining smaller by 1122 orders of magnitude compared to that of KM3NeT. This is because $P_{s\mu}$ depends on $\sin^{2}(L/L_{\text{res}})$, and the neutrino propagation length in Earth (along the direction of the 220 PeV neutrino detected by KM3NeT) is about 10 times shorter at IceCube. Therefore, at resonance, $P_{s\mu}$ is always expected to be at least two orders of magnitude larger for KM3NeT. The figure can be extended to larger values of $\epsilon_{ss}$ that would require heavier sterile neutrinos to induce a resonance. However, increasing the interaction strength would create tension with the data Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm . As argued in Harnik:2012ni , for a leptophobic $U(1)_{B}$ gauge boson that couples only to baryon number, many bounds are not applicable or they weaken. For example, with the leading production mechanism being through radiation off nuclei, stellar cooling bounds would be suppressed Harnik:2012ni . Further, supernova bounds for this scenario go only up-to light vector boson masses of $\sim 0.2$ GeV Rrapaj:2015wgs . Some of the most stringent constraints on such a model arise from anomaly cancellation Dror:2017ehi , with constraints for the vector boson coupling to baryons at $g_{b}\lesssim 1.85\times 10^{-3}$, for a benchmark mass of $\sim 0.5$ GeV. The coupling of the vector boson with the sterile sector can be large. If we allow a maximum coupling of $g_{b}^{\prime}\approx 1$ (remaining well within the perturbative limit), we obtain a maximum allowed value of $|\epsilon_{ss}|\lesssim 450$. Note that this completely rules out the red region in Fig. 2, preferred for a larger $\nu_{s}\to\nu_{\mu}$ oscillation probability at IceCube.
In Fig. 3, we show $P_{s\mu}$ as a function of $\epsilon_{ss}$ for both KM3NeT and IceCube and for a fixed sterile neutrino mass of $m_{s}=3$ keV. The resonance is obtained for $\epsilon_{ss}\simeq 150$, a parameter point where $P_{s\mu}$ for KM3NeT is two orders of magnitude larger than for IceCube. This difference in oscillation probability, i.e., in the generated muon neutrino flux, alleviates the tension between the two experiments by compensating for KM3NeT’s smaller effective area and shorter data-taking period. In Fig. 3, we also show the dotted line, which represents the probability for sterile-to-active oscillations in vacuum, i.e. the muon neutrino flux generated between the source and the Earth relative to the sterile neutrino flux. The muon neutrino flux generated in vacuum is subdominant compared to the one generated through resonance, as required, since otherwise the event rate would be dominated by neutrinos produced outside the Earth, which would not explain the tension.
@@
Figure 4: The sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the scenario with off-diagonal non-standard neutrino interactions is shown for both KM3NeT and IceCube for $m_{s}=500$ eV.
In our scenario, a sterile neutrino with $E_{\nu}=220$ PeV would mix with the active neutrino sector due to matter effects parameterized by $\epsilon_{\mu s}$. The effective mixing angle in matter in the limit of a tiny vacuum mixing angle $\theta$, large $m_{s}$, and $V_{\rm NC}=-V_{\rm CC}/2$ takes the form
@@
In the limit of $V_{\text{CC}}\gg m_{s}^{2}/E_{\nu}$, this further reduces to $\theta_{m}=(1/2)\tan^{-1}(4\epsilon_{\mu s})$. The effective mass squared difference in matter for $\theta\approx 0$ and $V_{\rm NC}=-V_{\rm CC}/2$ reads
@@
where $L$ is neutrino propagation length in Earth. In Fig. 4, we show the conversion probabilities for KM3NeT and IceCube as a function of $\epsilon_{\mu s}$ for $m_{s}=500\text{ eV}$ and $\theta=10^{-4}$. We have scrutinized the impact of the considered interaction on $P_{ee}$, $P_{\mu e}$, and $P_{\mu\mu}$ at low energies and found that $\epsilon_{\mu s}$ values up to $\mathcal{O}(1-10)$ have negligible effect on neutrino oscillations for $m_{s}\gtrsim 100$ eV. Hence, in the figure, we present oscillation probabilities for $\epsilon_{\mu s}$ values up to $10$. The sterile-to-active oscillation probability for KM3NeT is larger than that for IceCube by roughly two orders of magnitude for all presented values of $\epsilon_{\mu s}$. This can be understood from the oscillation formula in Eq. 8, which features an $L^{2}$ dependence when ${m_{\text{eff}}^{2}L}/{(4E_{\nu})}\lesssim\mathcal{O}(1)$, which is satisfied. Thus, the difference in the propagation length between KM3NeT and IceCube leads to the observed difference in the value of $P_{s\mu}$. While the ratio of $P_{s\mu}$ for KM3NeT and IceCube does not depend on $\epsilon_{\mu s}$, it should be noted that solutions with larger $\epsilon_{\mu s}$ are more favorable, as they lead to a larger generated active neutrino flux at KM3NeT, and thus require smaller incoming flux of sterile neutrinos. As in our previous scenario with resonance, the larger muon neutrino flux generated en route to KM3NeT alleviates the tension arising from the non-observation of $\mathcal{O}(100)$ PeV neutrinos in IceCube. For the energy dependence of the non-resonant scenario, see the Supplemental Material.
Note that, both the scenarios presented in our letter ensure that only sterile neutrinos at a particular energy range (and beyond) oscillate into active neutrinos in the presence of matter, so it is allowed for this individual transient event to have a much broader energy range in the incoming sterile flux. Namely, $\mathcal{O}(100)$ PeV neutrinos can oscillate into active neutrinos for the resonant solution. Whereas, for the non-standard interaction solution, for $E_{\nu}V_{\text{CC}}\gg m_{s}^{2}$ the sterile to active conversion probability would be enhanced significantly in the presence of matter effects; higher energies would be suppressed by an expected power law drop in the sterile neutrino flux.
In our study, we consider a transient sterile neutrino source; the required luminosity to observe a single event at KM3NeT depends on the emission duration from the source, the conversion probability $P_{s\mu}$ at Earth, and the source distance. For a purely sterile transient source, at a distance of $100$ Mpc, lasting $\sim\mathcal{O}(10^{3})$ s during the optimal observation window, the required time-averaged sterile neutrino luminosity in the $\mathcal{O}(150-300)$ PeV band would need to be $\sim\mathcal{O}(10^{50}-10^{51})$ erg/s. Larger values of $m_{s}$ and $\epsilon_{ss}$ (or $\epsilon_{\mu s}$), satisfying either the resonance condition or those corresponding to the non-standard interaction scenario, can lead to an enhanced $P_{s\mu}$ which can reduce the required luminosity by a factor of $\sim\mathcal{O}(10-100)$.